mathslover (mathslover):

Represent \(\large{\sqrt{5.6}}\) on the no. line

mathslover (mathslover):

Recently i got to know about how to represent \(\sqrt{2}\) on the no. line but now it is \(\sqrt{5.6}\)

OpenStudy (ash2326):

@mathslover 2.4 *2.4=5.76 2.3*2.3=5.29 so it will lie in between 2.3 and 2.4

mathslover (mathslover):

yes but how to exactly mark it on no. line @ash2326 ?

mathslover (mathslover):

"still reminding :("

OpenStudy (ash2326):

We have 5.29 5.6 and 5.76 5.6 is near to 5.76 so \(\sqrt {5.6}\) would be close to 2.4

mathslover (mathslover):

wht abt this?

OpenStudy (ash2326):

@mathslover if you need to represent it on a graph, then you need to find the exact value. Do you know how to find square root?

mathslover (mathslover):

yes @ash2326 but to find the exact value of AN IRRATIONAL NUMBER is impossible .. and i have to do it by only given manner that i showed But i do think that i am doing mistake somewhere ? @amistre64 @TuringTest @ash2326 @Vaidehi09 ..

OpenStudy (ash2326):

oops I meant a numeric value

mathslover (mathslover):

Yes but i dont know why the teacher is forcing me to do this only by above method tomorrow is my last exam please help .. this is the only topic where m having confusion

OpenStudy (ash2326):

Give me some time, I'll guide you

mathslover (mathslover):

ok take ur time .. sir!!!

OpenStudy (ash2326):

got it:) Let's start now

OpenStudy (ash2326):

We have to represent \(\sqrt {5.6}\) on number line First step draw a line say AB of 5.6 cm length

OpenStudy (ash2326):

@mathslover ????

mathslover (mathslover):

sorry for late reply @ash2326 ... this is what i did ? m i right ?

mathslover (mathslover):

@TuringTest can u check it out?

OpenStudy (ash2326):

Yeah you're right:D

mathslover (mathslover):

Oh k thanks a lot @ash2326

OpenStudy (zepp):

@mathslover Autodesk Sketchbook? ;D

mathslover (mathslover):

No @zepp not autodesk ..

OpenStudy (zepp):

Main colors looks quite similar